# SpectralEarth: Training Hyperspectral Foundation Models at Scale

### v1

@@ 5.2 Downstream Tasks
We append a linear layer to the pretrained encoders and train the models with Binary Cross-Entropy loss for 100 epochs.
Semantic segmentation on EnMAP-CDL/-NLCD and DESIS-CDL.
@@
We use a lightweight transfer protocol for the segmentation tasks, following previous works [26, 22]. We train for 100 epochs with a cross-entropy loss. The pretrained encoders are converted into segmentation models as follows:
@@
Parameter regression on Hyperview.
We use the training set of the Hyperview challenge [49] to evaluate cross-sensor transferability. Hyperview is a soil parameter estimation dataset comprising 1,732 hyperspectral patches with 150 spectral bands. Each patch is annotated with four soil parameters, including potassium ($K$), phosphorus pentoxide ($P_{2}O_{5}$), magnesium ($Mg$), and acidity $pH$. For this downstream task, we append two fully connected layers to the backbones. The models are optimized based on the normalized MSE criterion defined in Section 5.4 with the AdamW optimizer for 200 epochs.
5.3 Evaluation Protocols
We consider multiple evaluation protocols with various levels of fine-tuning of the pretrained models.
@@ 5.4 Evaluation Metrics
  • Segmentation: The mean Intersection over Union (mIoU) is reported for the EnMAP-CDL, EnMAP-NLCD and DESIS-CDL datasets.
@@
6 Benchmark Results
@@
Our main results are summarized in Tables 2 and 3. We include random initialization baselines for each downstream task, along with our pretrained models. Due to hundreds of spectral channels of the SpectralEarth patches, a comparison with RGB-ImageNet weights or multispectral pretrained models is of little use.
Evaluation protocol (##\## of trainable params.) Init Weights EnMAP-CORINE (F1 Score) EnMAP-CDL (mIoU) EnMAP-NLCD (mIoU)
Spec. RN50 Spec. ViT-S Spec. RN50 Spec. ViT-S Spec. RN50 Spec. ViT-S
Frozen Encoder (0) Random $70.53$ $70.42$ $44.72$ $46.52$ $35.86$ $36.85$
MoCo-V2 $73.97$ $73.60$ $\mathbf{51.66}$ $50.37$ $\mathbf{41.98}$ $39.85$
DINO $\mathbf{76.64}$ $\mathbf{75.06}$ $51.53$ $51.01$ $41.77$ $40.31$
MAE -- $72.72$- $\mathbf{51.37}$- $\mathbf{41.17}$
Full Fine-tuning ($>$20M) Random $78.31$ $77.78$ $57.53$ $55.07$ $\mathbf{48.18}$ $45.95$
MoCo-V2 $\mathbf{78.57}$ $78.40$ $\mathbf{58.10}$ $55.84$ $48.09$ $45.78$
DINO $77.98$ $78.34$ $57.77$ $55.70$ $47.75$ $45.71$
MAE -- $\mathbf{78.66}$- $\mathbf{57.66}$- $\mathbf{47.82}$
Fine-tune Adapter (56K) MoCo-V2 $76.27$ $76.12$ $\mathbf{55.36}$ $54.37$ $\mathbf{44.66}$ $43.40$
DINO $\mathbf{78.43}$ $\mathbf{77.95}$ $55.26$ $53.50$ $44.41$ $43.00$
MAE -- $76.80$- $\mathbf{54.61}$- $\mathbf{43.92}$
@@
Table 2: Benchmark results of the SpectralEarth downstream tasks for Spec. RN50 and Spec. ViT-S under different evaluation protocols: frozen backbone, full fine-tuning, and fine-tuning the spectral adapter only. We report the number of trainable parameters in the pretrained backbones, excluding task-specific heads. The best scores of each evaluation protocol-backbone are in bold.
Evaluation Protocol EnMAP-CORINE (F1 Score) EnMAP-CDL (mIoU) EnMAP-NLCD (mIoU)
B L H g B L H g B L H g
Training from Scratch $76.99$ $77.24$ $\mathbf{77.28}$ $76.85$ $54.79$ $54.50$ $\mathbf{54.83}$ $54.74$ $\mathbf{45.96}$ $45.62$ $45.53$ $45.58$
Frozen Encoder $74.72$ $75.07$ $\mathbf{76.06}$ $75.33$ $51.20$ $53.14$ $\mathbf{53.19}$ $52.77$ $40.52$ $42.88$ $\mathbf{43.32}$ $42.63$
Full Fine-tuning $79.05$ $79.18$ $\mathbf{79.80}$ $78.38$ $57.70$ $\mathbf{58.19}$ $58.06$ $57.86$ $48.10$ $\mathbf{48.37}$ $48.28$ $48.08$
Fine-tune Adapter $77.09$ $77.79$ $\mathbf{77.94}$ $77.86$ $54.79$ $\mathbf{55.35}$ $55.14$ $54.90$ $43.97$ $44.46$ $44.67$ $\mathbf{44.73}$
Table 3: Benchmark results for large vision transformers with varying sizes pre-trained on SpectralEarth with MAE. Models are evaluated under three protocols: frozen encoder, full fine-tuning, and fine-tuned adapter. Training from scratch is reported as a baseline. The best score for each protocol is highlighted in bold.
6.1 Comparing SSL Algorithms
@@
Table 2 summarizes our results for three SSL algorithms with the Spec. RN50 and Spec. ViT-S backbones.
EnMAP-CORINE.
@@
Pretrained models outperform random weights in linear evaluation for MoCo-V2 and DINO, with DINO yielding the best results, demonstrating the ability of joint-embedding methods to learn robust representations, even in the absence of fine-tuning. Additionally, we observe that MAE under-performs in linear probing, which is consistent with existing literature [41]. This can be attributed to the reconstruction pre-training objective, which focuses on low-level features, whereas joint-embedding methods learn more global and semantic representations. For full fine-tuning, pretrained models are on par with/slightly outperform training from scratch.
@@
Fine-tuning the adapter yields a notable performance boost compared to linear evaluation. This significant performance improvement aligns with reports in the literature [40]: fine-tuning early network layers is an efficient compromise between frozen weights vs. complete fine-tuning. In terms of F1-score, only fine-tuning the adapter ($0.3\%$ of network parameters) with DINO matches training from scratch, highlighting the efficient adaptability of our pretrained models.
EnMAP-CDL.
SSL pre-training yields improved results over random initialization, with MoCo overall performing best for Spec. RN-50, while MAE performs best for Spec. ViT-S with full fine-tuning. Moreover, fine-tuning the adapter enables the spectral convolutions to adjust to the distribution of the downstream dataset, effectively bridging the performance gap between linear evaluation and full fine-tuning. For the Spec. ViT-S, fine-tuning the spectral adapter is only half a point lower than training from scratch. As a general observation, Spec. RN50 slightly outperforms Spec. ViT-S on EnMAP-CDL for all evaluation protocols. This shows that, despite the rise of vision transformers, convolutional networks remain highly effective in remote sensing, especially with smaller datasets.
EnMAP-NLCD.
On the NLCD downstream task, the frozen pretrained backbones yield improvements over random weights but remain significantly worse than training from scratch. Notably, MAE outperforms training from scratch by ${}2\%$ of mIoU, demonstrating its strong transferability with fine-tuning. Conversely, training from scratch shows competitive performance, matching full fine-tuning for MoCo-V2 and DINO. This observation is aligned with findings from previous studies [25, 26] suggesting that training from scratch is a strong baseline. A discussion on the influence of the number of fine-tuning epochs is provided in Section 6.5.
6.2 Scaling-up Pretrained Models
Recent studies have shown significant performance gains through model scaling in SSL [27, 51]. Consequently, recent geospatial foundation models are exploring ViT architectures with hundreds of millions of parameters [63, 31, 34]. We investigate model scaling on SpectralEarth by pretraining Spec. ViT architectures of varying sizes: Base (B), Large (L), Huge (H), and Giant (g). Our findings are presented in Table 3. We observe that MAE consistently outperforms training from scratch across all model sizes. Remarkably, fine-tuning the adapter, which corresponds to 56K trainable parameters, matches or surpasses training from scratch for the EnMAP-CORINE and EnMAP-CDL tasks. Moreover, we observe initial benefits from scaling up the models up to Spec. ViT-L/H. However, additional increases in model size do not improve performance, possibly due to the size of the pretraining dataset. While SpectralEarth is significantly larger than previous HSI datasets, it remains smaller than the datasets used in computer vision for training state-of-the-art models [51]. This highlights the importance of scaling the data when training large vision transformers in remote sensing.
6.3 Efficient Fine-tuning
Model convergence.

### v2

@@ V-B Downstream Tasks
We append a linear layer to the pretrained encoders and train the models with Binary Cross-Entropy loss for 100 epochs with a batch size of 128. AdamW optimizer was employed with a cosine annealing scheduler. Data augmentation included random resized crop, horizontal and vertical flipping.
Semantic segmentation
@@
We use a lightweight transfer protocol for the segmentation tasks, following previous works [19, 40]. The pretrained encoders are converted into segmentation models as follows:
@@
We train for 100 epochs with a cross-entropy loss. AdamW optimizer was used with a batch size of 64 for EnMAP-NLCD/TreeMap and 32 for EnMAP-CDL/BDForet/EuroCrops and DESIS-CDL. Data augmentation included random resized crop, horizontal and vertical flipping.
Parameter regression - Hyperview
We use the dataset from the Hyperview challenge [29] to evaluate cross-sensor transferability. Hyperview is a soil parameter estimation dataset comprising 1,732 hyperspectral patches for training and 1,154 for testing, each with 150 spectral bands. Each patch is annotated with four soil parameters, including potassium ($K$), phosphorus pentoxide ($P_{2}O_{5}$), magnesium ($Mg$), and acidity $pH$. We resized all images in the Hyperview dataset to $\times128$ pixels. For this downstream task, we append two fully connected layers to the backbones. The models are optimized based on the normalized MSE criterion defined in Section V-D with the AdamW optimizer for 400 epochs and a batch size of 32. As for the other tasks, we used random resized crop, horizontal and vertical flipping during training.
Pixel-wise regression - HyBiomass
We evaluate our models on the above ground forest biomass dataset HyBiomass [86], which is a pixel regression task. The dataset consists of more than 34,000 patches distributed globally over forest areas. The above-ground biomass estimates are derived from the Global Ecosystem Dynamics Investigation (GEDI) lidar mission. We set the batch size to 128 and use the AdamW optimizer for 50 epochs. We use the same data augmentations as for the segmentation downstream tasks.
V-C Evaluation Protocols
We consider multiple evaluation protocols with various levels of fine-tuning of the pretrained models:
@@ V-D Evaluation Metrics
  • Semantic Segmentation: The mean Intersection over Union (mIoU) is reported for all semantic segmentation datasets.
@@
  • Pixel-wise Regression: we report the R² value for the HyBiomass dataset.
V-E Comparison Methods
We compare our pretrained models against several baselines and recent foundation models for hyperspectral imagery:
  • Random Frozen Encoder: We use this baseline as a sanity check to assess whether the learned representations provide useful features for downstream tasks.
  • Training from Scratch: A strong baseline that has shown competitive performance when trained with sufficient data and epochs [88].
  • DOFA-(B/L) [70]: A foundation model capable of adapting to various sensors given the input wavelengths. The model includes EnMAP data in its pretraining corpus. We evaluate both the base (DOFA-B) and large (DOFA-L) variants.
  • SpatSigma-(B/L) [16]: HyperSigma is a recent foundation model for hyperspectral imagery pretrained on Gaofen and Hyperion EO-1 data. We use only the spatial encoder for simplicity, as the original paper reports limited benefits when combining the spatial and spectral branches. We report results for the base and large variants and refer to them as SpatSigma-B and SpatSigma-L respectively. To mitigate the mismatch in the number of spectral bands, we resample the patch embedding layers with linear interpolation for our experiments.
VI Benchmark Results
@@
Our main results are summarized in Tables II and III. We include random initialization baselines for each downstream task, along with our pretrained models. Due to hundreds of spectral channels of the SpectralEarth patches, a comparison with RGB-ImageNet weights or multispectral pretrained models is of little use.
Model Configuration Classif. (F1) Segmentation (mIoU) Pix Reg. (R²)
Arch. Weights Protocol CORINE CDL NLCD EUROCROPS TREEMAP BDFORET BNETD HYBIOMASS
Random Frozen 69.92 $\pm$ 0.51 62.65 $\pm$ 0.80 35.94 $\pm$ 0.38 54.83 $\pm$ 0.83 35.87 $\pm$ 0.23 57.11 $\pm$ 0.80 41.35 $\pm$ 0.30 0.30 $\pm$ 0.01
Full FT 77.77 $\pm$ 0.38 77.28 $\pm$ 0.13 47.95 $\pm$ 0.04 68.93 $\pm$ 1.41 46.47 $\pm$ 0.04 75.50 $\pm$ 0.42 48.97 $\pm$ 0.10 0.48 $\pm$ 0.00
Spec. RN50 MoCo-V2 Frozen 74.07 $\pm$ 0.14 70.97 $\pm$ 0.19 41.87 $\pm$ 0.06 61.70 $\pm$ 0.08 41.67 $\pm$ 0.05 67.60 $\pm$ 0.34 44.84 $\pm$ 0.05 0.39 $\pm$ 0.00
Full FT 77.43 $\pm$ 0.46 77.53 $\pm$ 0.25 48.11 $\pm$ 0.03 70.95 $\pm$ 0.63 46.94 $\pm$ 0.03 78.56 $\pm$ 0.32 49.54 $\pm$ 0.04 0.50 $\pm$ 0.00
Adapter 76.82 $\pm$ 0.39 75.84 $\pm$ 0.11 44.62 $\pm$ 0.03 66.71 $\pm$ 0.23 44.84 $\pm$ 0.04 73.86 $\pm$ 0.19 47.46 $\pm$ 0.04 0.45 $\pm$ 0.00
DINO Frozen 76.61 $\pm$ 0.31 71.02 $\pm$ 0.13 41.83 $\pm$ 0.03 60.29 $\pm$ 0.29 42.03 $\pm$ 0.03 67.07 $\pm$ 0.11 44.32 $\pm$ 0.02 0.40 $\pm$ 0.00
Full FT 77.71 $\pm$ 0.55 77.49 $\pm$ 0.23 47.62 $\pm$ 0.07 69.52 $\pm$ 1.00 47.08 $\pm$ 0.13 74.68 $\pm$ 0.90 49.74 $\pm$ 0.31 0.50 $\pm$ 0.00
Adapter 76.96 $\pm$ 0.37 75.51 $\pm$ 0.15 44.38 $\pm$ 0.03 66.62 $\pm$ 0.35 44.70 $\pm$ 0.07 73.90 $\pm$ 0.11 47.21 $\pm$ 0.07 0.45 $\pm$ 0.00
Random Frozen 70.59 $\pm$ 0.62 65.96 $\pm$ 0.72 36.94 $\pm$ 0.30 55.78 $\pm$ 0.52 38.72 $\pm$ 0.25 61.89 $\pm$ 0.85 41.55 $\pm$ 0.22 0.34 $\pm$ 0.00
Full FT 77.63 $\pm$ 0.37 74.77 $\pm$ 0.22 45.31 $\pm$ 0.62 66.87 $\pm$ 0.62 44.43 $\pm$ 0.13 74.00 $\pm$ 0.49 46.57 $\pm$ 0.11 0.40 $\pm$ 0.02
MoCo-V2 Frozen 73.99 $\pm$ 0.04 71.40 $\pm$ 0.32 39.99 $\pm$ 0.08 60.21 $\pm$ 0.39 41.76 $\pm$ 0.06 68.50 $\pm$ 0.12 43.50 $\pm$ 0.05 0.40 $\pm$ 0.00
Full FT 78.37 $\pm$ 0.34 75.59 $\pm$ 0.14 45.92 $\pm$ 0.05 67.27 $\pm$ 0.29 45.03 $\pm$ 0.10 75.24 $\pm$ 0.28 47.02 $\pm$ 0.07 0.43 $\pm$ 0.01
Adapter 76.72 $\pm$ 0.35 75.28 $\pm$ 0.19 43.18 $\pm$ 0.08 66.46 $\pm$ 0.21 44.09 $\pm$ 0.07 73.85 $\pm$ 0.23 45.73 $\pm$ 0.10 0.44 $\pm$ 0.00
Spec. ViT-S DINO Frozen 75.11 $\pm$ 0.13 71.56 $\pm$ 0.35 40.47 $\pm$ 0.06 58.76 $\pm$ 0.27 41.89 $\pm$ 0.13 68.16 $\pm$ 0.43 43.12 $\pm$ 0.10 0.40 $\pm$ 0.00
Full FT 78.67 $\pm$ 0.34 76.10 $\pm$ 0.31 45.64 $\pm$ 0.15 65.56 $\pm$ 0.76 45.44 $\pm$ 0.15 73.56 $\pm$ 0.41 47.11 $\pm$ 0.14 0.48 $\pm$ 0.00
Adapter 77.00 $\pm$ 0.50 74.83 $\pm$ 0.07 42.86 $\pm$ 0.08 64.78 $\pm$ 0.65 43.54 $\pm$ 0.11 72.38 $\pm$ 0.71 44.99 $\pm$ 0.17 0.44 $\pm$ 0.00
MAE Frozen 72.85 $\pm$ 0.20 72.01 $\pm$ 0.28 41.08 $\pm$ 0.05 62.14 $\pm$ 0.18 41.19 $\pm$ 0.15 68.92 $\pm$ 0.73 43.34 $\pm$ 0.26 0.42 $\pm$ 0.00
Full FT 79.17 $\pm$ 0.55 77.16 $\pm$ 0.31 47.61 $\pm$ 0.10 69.87 $\pm$ 0.16 45.80 $\pm$ 0.15 76.54 $\pm$ 0.23 49.36 $\pm$ 0.07 0.51 $\pm$ 0.01
Adapter 76.75 $\pm$ 0.35 75.21 $\pm$ 0.21 43.64 $\pm$ 0.07 67.18 $\pm$ 0.58 43.45 $\pm$ 0.14 74.89 $\pm$ 0.47 45.89 $\pm$ 0.12 0.45 $\pm$ 0.01
@@
TABLE II: Benchmark results of the SpectralEarth downstream tasks for Spec. RN50 and Spec. ViT-S under different evaluation protocols: frozen backbone, full fine-tuning, and fine-tuning the spectral adapter only. The best score for each dataset-backbone pair is in bold, and the second best is underlined.
VI-A Comparing SSL Algorithms
@@
Table II summarizes our results for three SSL algorithms with the Spec. RN50 and Spec. ViT-S backbones.
Multi-label classification
@@
Pretrained models outperform random weights in linear evaluation for MoCo-V2 and DINO on EnMAP-CORINE, with DINO yielding the best results, demonstrating the ability of joint-embedding methods to learn robust representations, even in the absence of fine-tuning.
Additionally, we observe that MAE under-performs in linear probing, which is consistent with existing literature [89]. This can be attributed to the reconstruction pre-training objective, which focuses on low-level features, whereas joint-embedding methods learn more global and semantic representations. For full fine-tuning, CNN pretrained models with MoCo and DINO are on par with/slightly worse than training from scratch. In contrast, the Spec. ViT-S model benefits from pre-training in all settings compared to random initialization, with MAE yielding the best results.
@@
Fine-tuning the spectral adapter yields a notable performance boost compared to linear evaluation. This observation aligns with reports in the literature [90]: fine-tuning early network layers is an efficient compromise between frozen weights vs. complete fine-tuning. In terms of F1 score, only fine-tuning the adapter ($0.3\%$ of network parameters) is, in most settings, less than one point below training from scratch, highlighting the efficient adaptability of our pretrained models.
Semantic segmentation
Pretrained models with a frozen encoder consistently outperform random weights across all segmentation tasks. This shows that, without any fine-tuning, the models have learned useful features for segmentation. MoCo-V2 and DINO pre-trained models yield comparable results for both the CNN and the ViT backbones. However, MAE, as a feature extractor, consistently outperforms MoCo-V2 and DINO for Spec. ViT-S. This can be attributed to the pixel-level pre-training objective, which is more suitable for segmentation tasks (as opposed to image-level problems).
For complete fine-tuning, MoCo-V2 pre-trained CNNs consistently outperform training from scratch on all segmentation tasks, while DINO is on par with the random initialization baseline. Notably, the gaps are more pronounced for BDForet and EuroCrops, where MoCo-V2 yields an improvement of 3 and 2 points of mIoU, respectively. For the ViT model, MAE + fine-tuning performs best in all settings, yielding consistent improvements of 2 to 3 points of mIoU. Depending on the dataset, the second-best models are either MoCo-V2 or DINO, with smaller improvements compared to training from scratch.
Fine-tuning the spectral adapter significantly outperforms frozen encoder performance in all settings, showing the importance of tailoring the spectral feature extraction to the characteristics of the downstream task. Notably, fine-tuning the spectral adapter outperforms training from scratch on CDL, EuroCrops, and BD-Foret.
Compared to CNNs, ViTs perform slightly worse on segmentation problems. This can be attributed to multiple factors:
  1. 1.
    Training ViTs from scratch results in substantially lower performance than CNNs due to the lack of strong inductive biases. While pre-training significantly mitigates this gap, CNNs still retain a slight edge after fine-tuning;
  2. 2.
    The spatial patterns of the masks, which present high-frequency features partly due to label noise, are harder to learn with a ViT due to the fixed patch size compared to CNNs, which can preserve full-resolution feature maps. This effect can be better observed in the patch size experiment; see Figure 13.
Pixel-wise regression
The results on HyBiomass are in line with the segmentation downstream tasks. Our pre-trained frozen encoders significantly outperform the random encoder baseline for all SSL methods and backbones. Fine-tuning the spectral adapter yields a significant performance boost compared to the frozen encoder setting, and full fine-tuning works best. MoCo and DINO provide consistent improvements with fine-tuning over training from scratch for both the CNN and ViT backbones. The largest gain from pre-training is observed for Spec. ViT-S with MAE, which achieves an R² improvement of 0.11 compared to training from scratch. This configuration also yields the overall best result. Moreover, both the frozen encoder and the adapter-only fine-tuning outperform training from scratch for Spec. ViT-S with MAE. This is consistent with our segmentation results: without pre-trained weights, CNNs significantly outperform ViTs, which can be attributed to the lack of strong inductive biases in ViTs, requiring larger datasets to perform well.
Model Configuration Classif. (F1) Segmentation (mIoU) Pix Reg. (R²)
Arch. Protocol CORINE CDL NLCD EUROCROPS TREEMAP BDFORET BNETD HYBIOMASS
Spec. ViT Models
Spec. ViT-B Supervised 77.17 $\pm$ 0.72 74.65 $\pm$ 0.60 45.57 $\pm$ 0.14 66.14 $\pm$ 0.43 44.51 $\pm$ 0.13 73.55 $\pm$ 0.39 46.76 $\pm$ 0.10 0.37 $\pm$ 0.02
Frozen 75.04 $\pm$ 0.30 72.20 $\pm$ 0.16 40.82 $\pm$ 0.14 61.29 $\pm$ 0.22 39.74 $\pm$ 0.27 68.39 $\pm$ 0.20 43.37 $\pm$ 0.10 0.38 $\pm$ 0.01
Full FT 79.49 $\pm$ 0.26 77.44 $\pm$ 0.26 47.59 $\pm$ 0.23 69.34 $\pm$ 0.43 46.10 $\pm$ 0.04 76.30 $\pm$ 0.52 49.46 $\pm$ 0.06 0.51 $\pm$ 0.00
Adapter 77.63 $\pm$ 0.25 75.30 $\pm$ 0.32 43.81 $\pm$ 0.04 67.31 $\pm$ 0.43 43.65 $\pm$ 0.18 74.51 $\pm$ 0.15 45.52 $\pm$ 0.11 0.44 $\pm$ 0.01
Spec. ViT-L Supervised 76.85 $\pm$ 0.67 74.28 $\pm$ 0.34 45.49 $\pm$ 0.13 65.66 $\pm$ 0.46 44.52 $\pm$ 0.08 72.93 $\pm$ 0.59 46.66 $\pm$ 0.04 0.38 $\pm$ 0.02
Frozen 73.98 $\pm$ 0.21 72.72 $\pm$ 0.23 42.67 $\pm$ 0.03 62.13 $\pm$ 0.18 42.62 $\pm$ 0.19 71.07 $\pm$ 0.31 43.96 $\pm$ 0.18 0.42 $\pm$ 0.00
Full FT 79.25 $\pm$ 0.65 77.41 $\pm$ 0.26 47.84 $\pm$ 0.03 68.73 $\pm$ 0.36 46.10 $\pm$ 0.17 76.26 $\pm$ 0.73 48.96 $\pm$ 0.39 0.49 $\pm$ 0.00
Adapter 77.24 $\pm$ 0.07 75.94 $\pm$ 0.13 44.15 $\pm$ 0.07 67.69 $\pm$ 0.32 44.09 $\pm$ 0.09 74.89 $\pm$ 0.26 46.11 $\pm$ 0.12 0.45 $\pm$ 0.00
Spec. ViT-H Supervised 77.18 $\pm$ 0.44 73.72 $\pm$ 0.47 45.33 $\pm$ 0.12 64.77 $\pm$ 0.48 44.45 $\pm$ 0.17 72.98 $\pm$ 0.45 46.62 $\pm$ 0.14 0.34 $\pm$ 0.05
Frozen 73.96 $\pm$ 0.30 71.68 $\pm$ 0.21 41.87 $\pm$ 0.13 61.94 $\pm$ 0.11 41.95 $\pm$ 0.17 68.90 $\pm$ 0.12 43.48 $\pm$ 0.14 0.38 $\pm$ 0.00
Full FT 79.51 $\pm$ 0.30 77.26 $\pm$ 0.18 47.85 $\pm$ 0.14 69.13 $\pm$ 0.49 46.05 $\pm$ 0.13 76.57 $\pm$ 0.35 49.38 $\pm$ 0.23 0.46 $\pm$ 0.01
Adapter 77.23 $\pm$ 0.32 75.34 $\pm$ 0.22 43.90 $\pm$ 0.08 66.51 $\pm$ 0.14 43.75 $\pm$ 0.06 74.13 $\pm$ 0.23 45.56 $\pm$ 0.09 0.45 $\pm$ 0.01
Spec. ViT-g Supervised 76.94 $\pm$ 0.61 73.44 $\pm$ 0.38 45.03 $\pm$ 0.04 64.40 $\pm$ 0.40 44.38 $\pm$ 0.11 72.83 $\pm$ 0.50 46.34 $\pm$ 0.07 0.39 $\pm$ 0.01
Frozen 72.50 $\pm$ 0.26 68.39 $\pm$ 0.20 39.68 $\pm$ 0.03 57.11 $\pm$ 0.13 40.21 $\pm$ 0.07 62.99 $\pm$ 0.09 41.94 $\pm$ 0.08 0.36 $\pm$ 0.00
Full FT 78.20 $\pm$ 0.66 76.14 $\pm$ 0.23 46.61 $\pm$ 0.17 66.75 $\pm$ 0.49 45.18 $\pm$ 0.12 73.47 $\pm$ 0.58 48.26 $\pm$ 0.06 0.35 $\pm$ 0.07
Adapter 77.26 $\pm$ 0.50 75.25 $\pm$ 0.24 43.17 $\pm$ 0.09 66.84 $\pm$ 0.29 43.69 $\pm$ 0.05 74.06 $\pm$ 0.47 45.10 $\pm$ 0.06 0.41 $\pm$ 0.01
SOTA Models
DOFA-B Frozen 67.82 $\pm$ 0.15 58.99 $\pm$ 0.13 33.47 $\pm$ 0.07 45.44 $\pm$ 0.51 34.60 $\pm$ 0.20 53.22 $\pm$ 0.43 37.04 $\pm$ 0.10 0.28 $\pm$ 0.00
Full FT 74.47 $\pm$ 0.66 69.11 $\pm$ 0.30 42.60 $\pm$ 0.06 56.99 $\pm$ 0.25 41.98 $\pm$ 0.05 64.90 $\pm$ 0.17 42.31 $\pm$ 0.06 0.40 $\pm$ 0.00
DOFA-L Frozen 67.48 $\pm$ 0.13 60.64 $\pm$ 0.27 33.42 $\pm$ 0.08 46.11 $\pm$ 0.17 34.67 $\pm$ 0.14 51.26 $\pm$ 0.37 37.37 $\pm$ 0.08 0.28 $\pm$ 0.00
Full FT 74.72 $\pm$ 0.32 69.40 $\pm$ 0.44 43.06 $\pm$ 0.07 56.54 $\pm$ 0.55 42.18 $\pm$ 0.17 64.80 $\pm$ 0.37 43.06 $\pm$ 0.05 0.41 $\pm$ 0.00
SpatSigma-B Frozen 68.16 $\pm$ 0.33 63.49 $\pm$ 0.09 36.02 $\pm$ 0.06 50.67 $\pm$ 0.36 36.99 $\pm$ 0.09 58.20 $\pm$ 0.26 38.62 $\pm$ 0.12 0.31 $\pm$ 0.00
Full FT 76.71 $\pm$ 0.32 70.63 $\pm$ 0.55 43.18 $\pm$ 0.07 58.32 $\pm$ 0.19 42.93 $\pm$ 0.09 69.27 $\pm$ 0.35 42.54 $\pm$ 0.06 0.41 $\pm$ 0.00
SpatSigma-L Frozen 68.37 $\pm$ 0.33 61.72 $\pm$ 0.22 34.42 $\pm$ 0.30 47.86 $\pm$ 0.28 35.70 $\pm$ 0.09 54.94 $\pm$ 0.31 37.85 $\pm$ 0.06 0.27 $\pm$ 0.03
Full FT 76.66 $\pm$ 0.29 70.29 $\pm$ 0.21 43.31 $\pm$ 0.07 56.83 $\pm$ 1.01 42.60 $\pm$ 0.10 67.98 $\pm$ 0.45 42.96 $\pm$ 0.08 0.42 $\pm$ 0.00
TABLE III: Benchmark results for large vision transformers pre-trained on SpectralEarth with MAE. Models are evaluated under three protocols: frozen encoder, full fine-tuning, and fine-tuned adapter. Training from scratch is reported as a baseline. State-of-the-art models DOFA and HyperSigma are reported for comparison. The best score for each dataset is highlighted in bold, and the second best in underlined.
VI-B Scaling-up Pretrained Models
Recent studies have shown significant performance gains through model scaling in SSL [22, 11]. Consequently, recent geospatial foundation models are exploring ViT architectures with hundreds of millions of parameters [16, 53, 73]. We investigated model scaling on SpectralEarth by pre-training Spec. ViT architectures of varying sizes: Base (B), Large (L), Huge (H), and Giant (g). Our findings are presented in Table III.
We observe that MAE consistently outperforms training from scratch across all model sizes and datasets. Notably, scaling from Spec. ViT-B to ViT-L improves frozen encoder performance on multiple downstream tasks, including NLCD, EuroCrops, TreeMap, BDForet, and HyBiomass. Fine-tuning the adapter — only 56K additional parameters — surpasses training from scratch on CORINE, CDL, EuroCrops, BDForet, and HyBiomass. While scaling provides limited performance gains for some downstream tasks up to Spec. ViT-H, we find that Spec. ViT-B offers the best trade-off between performance and computational cost. Further increases in model size, as seen for Spec. ViT-g, degrades performance, possibly due to the size of the pretraining dataset. Although SpectralEarth is significantly larger than previous HSI datasets, it remains smaller than the datasets typically used to train billion-parameter models in computer vision [11]. This trend can also be observed for models such as DOFA and HyperSigma, where the large variants provide limited benefits over the base backbones. Our models outperform both DOFA and HyperSigma across all tasks and protocols, which we attribute to the architectural design of our backbones that better preserves fine-grained spatial information and the large EnMAP corpus in SpectralEarth. In particular, HyperSigma was trained on Hyperion EO-1 and Gaofen-5B, making it less competitive in EnMAP downstream tasks.
VI-C Efficient Fine-tuning
Model convergence